Tohban Report 2014-08-27
From RHESSI Wiki
|Start Date:||20 Aug 2014|
|End Date:||27 Aug 2014|
|Next Tohban:||Hazel Bain|
|List all reports|
There was low / medium solar activity this week. There is one AR producing several C class flares still on the disk now, but it will rotate off the disk in the next few days.
How many GOES flares occurred?
Flares above B, C, M, X class were 10 46 5 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 2 29 4 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 0 0 0 0
There were RHESSI flares/GOES flares 223 / 61 over the time range 19-Aug-14 26-Aug-14
The SSR is emptying at the end of the daily pass set, with a maximum fill of 25%. (Note that this is even after we made the change to normal / vigorous from active vigorous late last week.)
Because we have memory to spare, and because David would like eclipse data to aid a new wave of TGF work, we will turn on eclipse data for both the front and rear segments. The next tohban should keep a close eye on the memory after this change takes effect.
The segmentation status has not changed, with 1,3,8, and 9 stably segmented and D6 segmented most of the time, but losing segmentation during/following SAA passages. This segmentation is expected to become more stable over time.
There were no data gaps.
Detector 9 was noisy last week, so the voltage was dropped by 20 steps on August 20. (See commanding list below.) This succeeded in eliminating the noise.
Similarly, Detector 2 became noisy starting on the 20th, so its HV was dropped by 20 steps on August 21. This also worked.
The recurring light leak issue has reared its ugly head. As usual, this produces fast rates in D2 and D3 (and to a lesser extent in some of the others, notably D5 and D9) at times when Earth glow leaks into the cryostat. This does not harm the detectors but cuts into the livetime. Because D2 is unsegmented and D3 is segmented, the problem shows up in the front fast rates for D2 and the rears for D3. To mitigate the problem, the D2 front fast threshold was raised from 0x30 to 0x38 and the D3 rear fast threshold was raised from 0x45 to 0x50 on August 26. (See command list below.) An initial look shows that this didn't fix the problem, but we should wait for a bit more data before determining if more steps should be taken.
On August 20 (DOY 232) at 20:57 UT, several changes were made:
- Changed to normal_vigorous decimation.
- Lowered the voltage on Detector 9 by 20 counts as discussed in today's meeting. Detector 9 is now at 3600 V.
- Replaced detector 9 with detector 3 in the average live time calculation. We now use detectors 1,1,3, and 3.
- Started a spin up, which spun the spacecraft up from 14.5 to 15.0 rpm.
On August 21 (DOY 233) at 20:41:14 UT, the HV on D2 was lowered by 20 counts to 4117V.
On August 26 (DOY 238) the following threshold changes were made:
- G2 front fast was raised from 0x30 to 0x38
- G3 rear fast was rasied from 0x45 to 0x50
On August 27 (DOY 239) the following changes were made, at 20:39 UT:
- The software HV cut off temperature limit was increased from 130 K to 135 K. (Side note: we think the yellow and red warning levels are probably at 10 and 5 degrees below the cutoff.)
- Also, we are now collecting both front and rear segment events during eclipse, as well as during the day of course.
Several other topics were discussed; please see individual emails and presentations for details. Topics included:
- Brian presented latest data on cryostat temperatures and the cryocooler condition.
- Richard presented an update to visibility-based imaging routines.
- The possibility of changing the spacecraft pointing was discussed, and it was decided not to do so at this time.
- Pascal presented work on RHESSI imaging using the 3rd harmonic.
- Säm presented spectrograms and images for recent flares. Spectrograms show emission >25 keV in all the detectors. Promising!
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||+/- 4 min for most of week, changed to full eclipse data today.|
|Night time data (rears)||+/- 4 min for most of week, changed to full eclipse data today.|
|Detector problems?||Noise in 2, 9, but resolved.|